OpenStudy (brooke..help00):

I will be fanning and medaling everyone who helps me on each individual problem! @wholock Solve 2(x + 1) = 2x + 5. All real numbers No solution 0 3

OpenStudy (otherworldly):

2x + 2 = 2x + 5 >>> 2x - 2x = 5 - 2

OpenStudy (skullpatrol):

No solution

OpenStudy (skullpatrol):

2x - 2x = 5 - 2 >>> 0 = 3 False